OpenStudy (anonymous):

What is the solution to the equation? 1 - Z/5 = 9 A. z = 50 B. z = 44 C. z = –2 D. z = – 40

OpenStudy (anonymous):

What do you think first.

OpenStudy (anonymous):

C

OpenStudy (anonymous):

now show working why

OpenStudy (anonymous):

idk I thought you were going to explain to me ;/

OpenStudy (anonymous):

you can rearrange the equation so that Z is left on one side. so first take the one to the other side, you have to -1 to both sides, this becomes. \[\frac{ -Z }{ 5 } = 9-1\] this becomes \[\frac{ -Z }{ 5 } = 8\] multiply both sides by 5 to move the 5 from the left hand side to the right \[-Z=8\times5\] \[-Z=40\] to move the negative, just multiply both sides by -1 \[Z=-40\] the answer is d

OpenStudy (anonymous):

thank you